[Brahms-l] Shift report 20050304 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Mar 04 2005 - 16:12:32 EST
Shifter leader was RD
Shifters were 


1. Only one store during this shift
2. collected run 13564 (while Angelika wagged the beam)
   run 13565 (with gap cleaning being done during run)
   run 13566 300K good trig 2

3. C405 HV channels found tripped. It can only hold -2600Volts

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 4 16:14:39 2005

This archive was generated by hypermail 2.1.8 : Fri Mar 04 2005 - 16:14:54 EST